OpenStudy (saifoo.khan):

What is the third–degree polynomial function such that f(0) = –18 and whose zeros are 1, 2, and 3? f(x) = 2x3 – 12x2 + 18x – 18 f(x) = 3x3 + 18x2 – 33x – 18 f(x) = 2x3 + 12x2 – 18x – 18 f(x) = 3x3 – 18x2 + 33x – 18

OpenStudy (ybarrap):

f(x) = 3x3 – 18x2 + 33x – 18

OpenStudy (saifoo.khan):

How?

OpenStudy (saifoo.khan):

are u sure that's the answer?

OpenStudy (saifoo.khan):

Oh i got it.. :D

OpenStudy (saifoo.khan):

thanks

OpenStudy (ybarrap):

By elimination: Using Descartes' Rule of Signs only 1st and 4th equation makes sense Plugging in the number 1, only 4th equation makes sense, eliminating 1st Or more directly: f(x)=3x3 – 18x2 + 33x – 18 =(x-1)(3x^2 - 15x + 18) = (x-1)3(x^2 - 5x + 6) =3(x-1)(x-2)(x-3) f(x) = 0 when x = 1,2 or 3

OpenStudy (ybarrap):

most factoring techniques I know weren't working so I started with (x-1) as common factor and everything worked out

OpenStudy (ybarrap):

I could have also used The Rational Root Theorem, which would have given me these rational roots. DRoS shows that roots are positive and real, another hint. Only factors of 6 are 1,2,3 and 6 as possible roots (based on RRT) of these 3 turned out to be the roots required.